# Oscillations Make Neural Networks Robust to Quantization

### v1

@@ Abstract
To see how this gives rise to oscillations, for an arbitrary $w$, denote $w_{0}$ the upper discretization threshold $w_{0}=q(w)+s/2$. For $\varepsilon\in(0,s/2)$ note that we have $q(w_{0}-\varepsilon)=q(w)$ and $q(w_{0}+\varepsilon)=q(w)+s$ so that
$\mathbf{\Delta}(w_{0}+\varepsilon)$ $=q(w)+s/2+\varepsilon-(q(w)+s)$ (15)
$=-s/2+\varepsilon,$ (16)
$\mathbf{\Delta}(w_{0}-\varepsilon)$ $=q(w)+s/2-\varepsilon-q(w)$ (17)
$=s/2-\varepsilon.$ (18)
@@
Assuming $x\neq 0$, the negative STE gradient “flips" from $-s/2$ to $s/2$ as the weight $w$ passes the quantization threshold $w_{0}$ from above, pushing the weight back towards the threshold. We note that the STE gradient is 00 at the special value $w=q(w)$, but the preceding argument shows that this is an unstable critical point and gradient noise will immediately cause the weights to move away from it. When combined with (stochastic) gradient descent and a finite discretization timestep we can identify this as the driving mechanism behind oscillations during training with QAT (Fig. 1).
@@
We can also see how the dynamics lead to clustering around quantization thresholds by looking at the sign of $\mathbf{\Delta}$ for different values of $w$. For a weight $w$ let $d_{\text{low}}(w)$ and $d_{\text{up}}(w)$ denote the distance from $w$ to the upper and lower thresholds, $d_{\text{low}}(w)=w-\bigl{(}q(w)-\frac{s}{2}\bigr{)}=\mathbf{\Delta}(w)+\tfrac{s}{2}$ and $d_{\text{up}}(w)=\bigl{(}q(w)+\frac{s}{2}\bigr{)}-w=\tfrac{s}{2}-\mathbf{\Delta}(w)$ respectively. If $w$ is closest to the upper threshold we have
$d_{\text{up}}<d_{\text{low}}\Longrightarrow\tfrac{s}{2}-\mathbf{\Delta}<\mathbf{\Delta}+\tfrac{s}{2}\Longrightarrow\mathbf{\Delta}>0$ (19)
@@
$d_{\text{low}}<d_{\text{up}}\Longrightarrow\mathbf{\Delta}+\tfrac{s}{2}<\tfrac{s}{2}-\mathbf{\Delta}\Longrightarrow\mathbf{\Delta}<0$ (20)
We emphasize that this mechanism causes the weights to move towards the quantization thresholds (the edges of quantization "bins") as opposed to the quantization levels (the centers of the quantization "bins").
Section 5 Regularization Method
Based on our theoretical observations in the one weight linear model, we now investigate empirically if the mechanism in Eq. (14) is sufficient to introduce weight oscillations in neural networks.
From the quantization difference in Eq. 10 and the STE gradient derived in Eq. 14, we have:
$\frac{\partial\mathcal{L}(q(w))}{\partial w}$ $=\frac{\partial\mathcal{L}(w)}{\partial w}-x^{2}\mathbf{\Delta}(w)$ (21)
@@
In order to emulate the effects of QAT, we propose a regularization term so that the training objective becomes:
$\mathcal{L}(q(\mathbf{w}))=\mathcal{L}(\mathbf{w})+\mathcal{R_{\lambda}}(\mathbf{w})$ (22)
@@
Here $\lambda\geq 0$ is a hyperparameter that controls the amount of regularization, $\ell$ ranges over the layers in the model and $i$ over the weights in each layer.
Using the STE, ^$\frac{\hat{\partial}q}{\partial\mathbf{w}}=1$, we have the following expression for the gradient:
^^$\frac{\hat{\partial}}{\hat{\partial}w^{\ell}_{i}}\mathcal{R_{\lambda}}(\mathbf{w})=\frac{\lambda}{n_{\ell}}\left(q(w^{\ell}_{i})-w^{\ell}_{i}\right)=-\frac{\lambda}{n_{\ell}}\mathbf{\Delta}(w_{i}^{\ell}).$ (24)
By the same reasoning as in Sec. 4 this pulls the weight $w_{i}^{\ell}$ towards the quantization threshold and causes the gradient to “flip" as $w_{i}^{\ell}$ crosses the threshold. We expect this to lead to oscillations based on the same mechanism as in the model from Sec. 4.
Figures 2 and  3 show the results of an experiment where we observe the weight distributions, and measured the oscillations, during training of a neural network (ResNet-18) with varying degrees of regularization, respectively. For comparison purposes the figures also shows the weight distributions and oscillations observed during training with QAT. Using the definition of an oscillation established in Nagel et al. (2022), we count an oscillation at epoch $i>1$ if $q(w_{t})\neq q(w_{t-1})$ and the direction of the change in the quantized space is opposite to that of the previous change. We note though that this method of counting misses the first threshold crossing during an oscillation A.5
Our first observation is that QAT displays more oscillations – also seen as clustering around the quantization threshold in Fig. 2-a) – than a baseline model without QAT or regularization (corresponding to $\lambda=0$ in Fig. 3-b)) . As we increase $\lambda$ we observe that the number of oscillations as well as the clustering increases. This confirms that our regularizer can indeed induce oscillations similar to QAT during the training of deep neural networks. At $\lambda=1$ (Fig. 3-c)) the number of oscillations observed with our regularizer is similar to the behaviour of QAT, lending support to our hypothesis that the mechanism in (15) is indeed at the root of the oscillations observed when training neural networks with QAT.
@@
Figure 3: The plots show the distribution of weights with oscillation counts $>0$ when training with a) QAT and b)-d) our regularizer for different values of $\lambda$. Here $\lambda=0$ corresponds to a full precision model where our regularizer has no influence on training. The y-axis represents the percentage of total weights in the first convolutional layer of a ResNet-18 trained from scratch for 50 epochs, while the x-axis shows the oscillation count. Following the oscillation definition from Nagel et al. (2022), we count oscillations at each epoch during training. The results demonstrate that QAT produces a significantly higher proportion of oscillating weights compared to $\lambda=0$. Furthermore, we observe that as we increase $\lambda$ a greater percentage of weights oscillates.
Section 6 Experiments & Results
In this section we empirically try to answer the question: is it sufficient to induce weight oscillations during training in order to get the benefits of QAT?
@@
We answer this question mostly affirmatively for ResNet and Vision Transformer architectures, based on the results of training ResNet-18 and Tiny ViT on the CIFAR-10 dataset. This is both in a training-from-scratch setting and when fine-tuning pretrained models. In all our experiments we use the regularizer $\mathcal{R}_{\lambda}$ defined in Eq. (23) to induce oscillations.
@@
In the following subsections we first describe the experimental setup, then we present the accuracy results from training-from-scratch and fine-tuning models trained with different quantization levels for the quantizer in $\mathcal{R}_{\lambda}$ or QAT and finally, we present the cross-bit accuracy of the fine-tuned models. We train models at ternary (3 possible values: -1, 0, 1), 3-bit and 4-bit. This is in line with contemporary research, where the emphasis lies on quantization at 4-bit and below since the challenges of maintaining accuracy are more significant compared to quantization at higher bit widths.
Section 6.1 Experimental setup
@@
We conducted our experiments using the CIFAR-10 dataset Krizhevsky et al. (2009) without data augmentation. We evaluated three architectures; A multi-layer perceptron with 5 hidden layers and 256 neurons per layer (MLP5), ResNet-18 He et al. (2016) and Tiny Vision transformer (Tiny ViT) Wu et al. (2022).
@@
For each architecture we used the Adam optimizer Kingma (2014) and tested multiple configurations: A baseline model to establish optimal floating-point accuracy and post-training quantization (PTQ) performance, a model with QAT and a model with our approach. The two latter configurations are trained using a ternary, 3-bit, and 4-bit quantizer.
@@
Training from Scratch For the MLP5 architecture, we used a learning rate of $10^{-3}$ and regularization parameter $\lambda$=1. The ResNet-18 was trained with a learning rate of $10^{-3}$ and $\lambda$=0.75 (see Appx. A.2 for our hyperparameter selection). We modified the ResNet-18 architecture by replacing the input layer with a smaller $3\times 3$ kernel and adapting the final layer for 10-class classification of both ResNet-18 and Tiny ViT. Training proceeded for a maximum of 100 epochs with early stopping triggered after 10 epochs without improvement in validation performance. For quantized models, we monitored the quantized validation accuracy at the target bit precision, while for the baseline, we tracked floating-point accuracy.
@@
Fine-tuning We fine-tuned two ImageNet-1k Deng et al. (2009) pre-trained models on CIFAR-10: a Tiny ViT (learning rate: $10^{-4}$, $\lambda$=1) and a ResNet-18 (learning rate: $10^{-3}$, $\lambda$=1). To maintain compatibility with the pre-trained architectures, we upsampled CIFAR-10 images to $224\times 224$ pixels. The $\lambda$ parameter selection process for Tiny ViT is detailed in Appx. A.2. Fine-tuning continued for up to 200 epochs, with early stopping after 30 epochs without improvement, using the same accuracy metrics as training from scratch.
@@
Quantization We implemented weight quantization using a per-tensor uniform symmetric quantizer as defined in Eq. 1. The quantization range was determined by computing minimum and maximum values per layer. In our implementation of ResNet-18 (11M parameters) all layers except batch normalization were quantized, covering 99.96% of parameters. For Tiny ViT (5.5M parameters) quantization was applied to MLP, Self-Attention, and key-query-value projection layers, encompassing 97.18% of parameters. And lastly for the MLP5 model all layers were quantized.
Section 6.2 Training-from-scratch
@@
Table 1 shows the results from training an MLP and ResNet-18 from scratch on the CIFAR-10 dataset. Our regularization method (OsciQuant) demonstrates improvements compared to the PTQ baseline from ternary quantization. More importantly, it also matches the performance of QAT at bit widths of 3 and 4.
@@
For both models we see that at 3-bit and 4-bit, our method exhibits similar performance as QAT but with less variability, while not differing significantly in the average number of training epochs required. With both models, QAT and OsciQuant are competitive with the full-precision baseline, although we observe an increased number of training epochs. Notably, both OsciQuant and QAT significantly outperform PTQ when applied to the full precision baseline.
Model Quantization method Accuracy Mean Epochs
MLP5 Baseline FP32 51.43 $\pm$ 0.39 14
Ternary PTQ 10.00 $\pm$ 0.02 14
Ternary QAT 49.20 $\pm$ 1.34 24
Ternary OsciQuant 36.49 $\pm$ 0.51 14
3-bit PTQ 20.97 $\pm$ 5.64 14
3-bit QAT 50.53 $\pm$ 1.43 33
3-bit OsciQuant 48.48 $\pm$ 0.29 15
4-bit PTQ 46.50 $\pm$ 0.76 14
4-bit QAT 51.39 $\pm$ 0.60 26
4-bit OsciQuant 50.72 $\pm$ 0.47 19
ResNet-18 Baseline FP32 83.26 $\pm$ 1.07 24
Ternary PTQ 10.00 $\pm$ 0.01 24
Ternary QAT 79.62 $\pm$ 6.42 42
Ternary OsciQuant 61.5 $\pm$ 1.82 56
3-bit PTQ 77.79 $\pm$ 4.0 24
3-bit QAT 82.51 $\pm$ 2.14 37
3-bit OsciQuant 81.77 $\pm$ 0.46 41
4-bit PTQ 82.11 $\pm$ 1.21 24
4-bit QAT 82.66 $\pm$ 2.57 28
4-bit OsciQuant 83.74 $\pm$ 0.59 32
Table 1: Comparison of accuracy when training from scratch on CIFAR-10. Results show classification accuracy and mean training epochs for MLP5 and ResNet-18 across different quantization approaches and bit-widths. Results is means and standard deviations over 5 random seeds.
Model Quantization method Accuracy Mean Epochs
ResNet-18 Baseline FP32 88.50 $\pm$ 0.64 4
Ternary PTQ 10.01 $\pm$ 0.01 4
Ternary QAT 77.02 $\pm$ 7.57 47
Ternary OsciQuant 44.59 $\pm$ 3.30 35
3-bit PTQ 10.28 $\pm$ 0.48 4
3-bit QAT 85.69 $\pm$ 1.83 25
3-bit OsciQuant 84.94 $\pm$ 1.59 27
4-bit PTQ 35.56 $\pm$ 9.05 4
4-bit QAT 87.71 $\pm$ 1.14 26
4-bit OsciQuant 87.08 $\pm$ 0.72 24
Tiny ViT Baseline FP32 96.11 $\pm$ 0.31 6
Ternary PTQ 9.39 $\pm$ 1.11 6
Ternary QAT 73.53 $\pm$ 0.77 140
Ternary OsciQuant 13.51 $\pm$ 1.32 28
3-bit PTQ 11.56 $\pm$ 1.99 6
3-bit QAT 88.13 $\pm$ 0.60 131
3-bit OsciQuant 88.68 $\pm$ 1.08 108
4-bit PTQ 21.57 $\pm$ 5.33 6
4-bit QAT 94.96 $\pm$ 0.33 57
4-bit OsciQuant 94.82 $\pm$ 0.51 90
@@
Table 2: Comparison of accuracy when fine-tuning on models pre-trained on ImageNet-1k. Results show classification accuracy and mean training epochs for MLP5 and ResNet-18 across different quantization approaches and bit-widths. Results is means and standard deviations over 5 random seeds.
Section 6.3 Fine-tuning
@@
Table 2 summarizes the test accuracies for fine-tuning using our OsciQuant method and QAT on ResNet-18 and Tiny ViT architectures. The observations are roughly in line with the results observed for training from scratch in the previous section with the exception of the number of epochs required for fine-tuning.
@@
On the ResNet architecture both QAT and our model train for significantly longer than the full precision baseline. As is the case for training from scratch, we see an increase in ternary performance compared to PTQ, but QAT still ourperforms our method in the ternary setting. Our regularization and QAT show comparable performance when quantized at 3 bits and 4 bits, while achieving test accuracy close to the full precision model at 4-bits.
The general trend regarding accuracy is identical for the vision transformer experiments, while we again note the high number of epochs require for both methods when fine-tuning, compared to the full precision baseline.
Section 6.4 Robustness to cross-bit quantization
As described above, the goal of our proposed regularization term is to train a model that maintains performance after quantization. Since the regularization term involves a quantization operator, we need to choose the quantization level in the regularization term. In this experiment we evaluated the robustness of our method and QAT towards quantization at levels different from the ones used during training.
For OsciQuant, we applied a regularization term with the training bit width during training and applied PTQ after training finished at a different quantization level. For QAT we trained using the training bit width and afterwards applied PTQ to the latent weights. For each method we also evaluated the corresponding model without PTQ, directly using the latent weights for inference (reported as FP32).
Table 3 shows the results from the experiment. A first observation is that the models produced by our method consistently achieve nearly full-precision accuracy when quantized at 8-bit or when used without quantization, irrespective of the quantization level used during training. This contrasts with QAT, which produces a viable 8-bit or full-precision model only when trained with at least 4-bit.
Furthermore we see that our method mostly maintains performance when trained at 3 or 4-bit and quantized at bit level of 3 or 4-bit. QAT also achieves this for Tiny ViT but for ResNet, the accuracy of QAT trained at 3-bit and quantized at other bit widths is barely above random guessing.
Regarding training with ternary quantization, we see that our method produces models that achieve near full precision performance for ResNet when quantized at 3-bit or higher. Ternary training for ViT is somewhat peculiar in that it fails to produce a model that is viable when quantized to ternary, whereas the performance of the resulting models starts to show a high level of variability at 4-bit and finally reaches close to full-precision accuracy at 8-bit. In contrast, for both ResNet and ViT, the performance of QAT degrades completely to random guessing when trained with ternary quantization and evaluated at any other quantization level.
@@
ResNet-18 Baseline (PTQ) 88.50 $\pm$ 0.64 10.01 $\pm$ 0.01 10.28 $\pm$ 0.48 35.56 $\pm$ 9.05 88.45 $\pm$ 0.64
Ternary QAT 10.39 $\pm$ 0.71 77.02 $\pm$ 7.57 9.75 $\pm$ 0.77 10.03 $\pm$ 0.51 10.35 $\pm$ 0.63
Ternary OsciQuant 87.44 $\pm$ 0.56 44.59 $\pm$ 3.30 85.42 $\pm$ 1.13 87.03 $\pm$ 0.65 87.42 $\pm$ 0.56
3-bit QAT 16.89 $\pm$ 4.97 10.01 $\pm$ 0.04 85.69 $\pm$ 1.83 17.42 $\pm$ 4.96 16.56 $\pm$ 4.32
3-bit OsciQuant 87.86 $\pm$ 0.42 20.19 $\pm$ 10.74 84.94 $\pm$ 1.59 87.56 $\pm$ 0.38 87.86 $\pm$ 0.42
4-bit QAT 87.75 $\pm$ 1.13 10.13 $\pm$ 0.29 82.08 $\pm$ 6.25 87.71 $\pm$ 1.14 87.76 $\pm$ 1.12
4-bit OsciQuant 87.85 $\pm$ 0.49 11.91 $\pm$ 0.87 85.57 $\pm$ 1.10 87.08 $\pm$ 0.72 87.87 $\pm$ 0.49
Tiny ViT Baseline (PTQ) 96.11 $\pm$ 0.31 9.39 $\pm$ 1.11 11.56 $\pm$ 1.99 21.57 $\pm$ 5.33 96.03 $\pm$ 0.34
Ternary QAT 10.62 $\pm$ 1.29 73.53 $\pm$ 0.77 11.52 $\pm$ 1.82 11.13 $\pm$ 1.75 10.61 $\pm$ 1.26
Ternary OsciQuant 95.79 $\pm$ 0.58 13.51 $\pm$ 1.32 12.53 $\pm$ 3.66 54.93 $\pm$ 27.32 95.76 $\pm$ 0.59
3-bit QAT 86.94 $\pm$ 0.91 19.78 $\pm$ 6.04 88.13 $\pm$ 0.60 86.69 $\pm$ 0.62 86.95 $\pm$ 0.89
3-bit OsciQuant 96.47 $\pm$ 0.11 9.48 $\pm$ 1.64 88.68 $\pm$ 1.08 95.35 $\pm$ 0.18 96.50 $\pm$ 0.11
4-bit QAT 95.14 $\pm$ 0.29 $1$$\pm$ 1.84 59.86 $\pm$ 19.95 94.96 $\pm$ 0.33 95.13 $\pm$ 0.28
4-bit OsciQuant 96.54 $\pm$ 0.09 11.90 $\pm$ 1.29 70.23 $\pm$ 12.75 94.82 $\pm$ 0.51 96.55 $\pm$ 0.09

### v2

@@ 3.2 Oscillation Mechanism
To see how the observations in Sec. 3.1 give rise to oscillations, let $w_{0}$ denote the upper quantization threshold for an arbitrary weight $w$, defined as $w_{0}=q(w)+s/2$. For $\varepsilon\in(0,s/2)$, note that we have $q(w_{0}-\varepsilon)=q(w)$ and $q(w_{0}+\varepsilon)=q(w)+s$, so that
$\mathbf{\Delta}(w_{0}+\varepsilon)$ $=q(w)+s/2+\varepsilon-(q(w)+s)=-s/2+\varepsilon,$ (13)
$\mathbf{\Delta}(w_{0}-\varepsilon)$ $=q(w)+s/2-\varepsilon-q(w)=s/2-\varepsilon.$ (14)
@@
Assuming $x\neq 0$, the negative STE gradient “flips” from s/2-s/2 to +s/2+s/2 as the weight $w$ passes the quantization threshold $w_{0}$, pushing the weight back towards the threshold. We note that the STE gradient is zero at the special value $w=q(w)$, but the preceding argument shows that this is an unstable critical point and gradient noise will immediately cause the weights to move away from it. When combined with (stochastic) gradient descent and a finite discretization timestep we can identify this as the driving mechanism behind oscillations during training with QAT (Fig. 1).
3.3 Weight Clustering
@@
We can also see how the dynamics described above can lead to weight clustering around quantization thresholds by looking at the sign of $\mathbf{\Delta}$ for different values of $w$. For a weight $w$ let $d_{\text{low}}(w)$ and $d_{\text{up}}(w)$ denote the distance from $w$ to the upper and lower thresholds, $d_{\text{low}}(w)=w-\bigl(q(w)-\frac{s}{2}\bigr)=\mathbf{\Delta}(w)+\tfrac{s}{2}$ and $d_{\text{up}}(w)=\bigl(q(w)+\frac{s}{2}\bigr)-w=\tfrac{s}{2}-\mathbf{\Delta}(w)$ respectively. If $w$ is closest to the upper threshold we have
$d_{\text{up}}<d_{\text{low}}\Longrightarrow\tfrac{s}{2}-\mathbf{\Delta}<\mathbf{\Delta}+\tfrac{s}{2}\Longrightarrow\mathbf{\Delta}>0.$ (15)
@@
$d_{\text{low}}<d_{\text{up}}\Longrightarrow\mathbf{\Delta}+\tfrac{s}{2}<\tfrac{s}{2}-\mathbf{\Delta}\Longrightarrow\mathbf{\Delta}<0.$ (16)
We emphasize that this mechanism causes the weight to move towards the quantization thresholds (the edges of quantization “bins”) as opposed to the quantization levels (the centers of the quantization “bins”). As shown above, the magnitude of the pull towards the threshold increases as the weight approaches it. The weight will therefore eventually cross the threshold and start oscillating, unless $L(w)$ and $\delta_{L}$ exactly cancel out, which is unlikely to happen with a finite gradient descent step size.
4 Regularization Method
Our theoretical observations in the linear model in Sec. 3, show that the oscillation component is the only part that differentiates QAT from standard full-precision training. We now confirm empirically that the mechanism in Eq. (12) is sufficient to introduce weight oscillations during training of neural networks, and study if this also results in QAT-like behavior with respect to the quantization noise.
From the quantization difference in Eq. (9) and the STE gradient derived in Eq. (12), we have
$\frac{\partial\mathcal{L}(q(w))}{\partial w}$ $=\frac{\partial\mathcal{L}(w)}{\partial w}-x^{2}\mathbf{\Delta}(w),$ (17)
@@ 4 Regularization Method
In order to emulate the effects described in Section 3, we propose a regularization term so that the training objective becomes
$\mathcal{L}(\mathbf{w})+\mathcal{R_{\lambda}}(\mathbf{w}),$ (18)
@@
Here $\lambda\geq 0$ is a hyperparameter that controls the amount of regularization, $\ell$ ranges over the layers in the model and $i$ over the weights in each layer. In this term, we replaced the factor $x^{2}$ by a hyperparameter $\lambda$, since the precise expression of $x^{2}$ is specific to the model studied in Sec. 3. We empirically find that this regularizer is sufficient to induce oscillations. The exploration of the design space of oscillation-inducing regularizers, including layer-dependent and/or adaptive scale factors, is left to future work.
Using the STE, $\frac{\hat{\partial}q}{\partial\mathbf{w}}=1$, we have the following expression for the gradient:
$\frac{\hat{\partial}}{\hat{\partial}w^{\ell}_{i}}\mathcal{R_{\lambda}}(\mathbf{w})=\frac{\lambda}{n_{\ell}}\left(q(w^{\ell}_{i})-w^{\ell}_{i}\right)=-\frac{\lambda}{n_{\ell}}\mathbf{\Delta}(w_{i}^{\ell}).$ (20)
By the same reasoning as in Sec. 3 this pulls the weight $w_{i}^{\ell}$ towards the quantization threshold and causes the gradient to “flip” as $w_{i}^{\ell}$ crosses the threshold. We expect this to lead to oscillations based on the same mechanism as in the model from Sec. 3.
Figures 2 and  3 show the results of an experiment where we observe the weight distributions, and measured the oscillations, during training of a neural network (ResNet-18) with varying degrees of regularization, respectively. For comparison, the figures also show the weight distributions and oscillations observed during training with QAT.
Comparison t-score p-value
$\lambda_{0}$ vs. $\lambda_{1}$ $25.13$ <0.001<0.001
$\lambda_{0}$ vs. $\lambda_{10}$ $38.86$ <0.001<0.001
$\lambda_{0}$ vs. QAT $18.83$ <0.001<0.001
Table 1: Welch’s t-test comparing the per-weight oscillation count after 50 epochs of training, for the different cases in Fig. 3. We note how the baseline $\lambda=0$ differs significantly from QAT and regularization induced oscillations with $\lambda=1,10$.
Our first observation is that QAT displays more oscillations Fig. 3-(a) than a baseline model without QAT or regularization (corresponding to $\lambda=0$ in Fig. 3-(b)). As we increase $\lambda$ we observe that the number of oscillations as well as the clustering increases. This confirms that the regularizer from Eq. (19) can indeed induce oscillations similar to QAT during the training of deep neural networks. At $\lambda=1$ (Fig. 3-(c)) the number of oscillations observed with regularization is similar to the behavior of QAT, lending support to our hypothesis that the mechanism in Eq. (13) is indeed at the root of the oscillations observed when training neural networks with QAT. We use Welch’s $t$-test to test if the oscillation counts in Fig. 3, which is the per-weight oscillation count after 50 epochs, are significantly different to the $\lambda=0$ baseline, reported in Table 1. Each of the pairwise comparison shows that the distribution of oscillations are significantly different.
@@
Figure 3: Distribution of weight oscillations. The plots show the distribution of weights with oscillation counts >0>0 when training with a) QAT and b)-d) the regularizer for different values of $\lambda$. Here $\lambda=0$ corresponds to a full-precision model where the regularizer has no influence on training. The y-axis represents the percentage of total weights in the first convolutional layer of a ResNet-18 trained from scratch for 50 epochs, while the x-axis shows the oscillation count after 50 epochs. Following the oscillation definition from nagel2022overcoming, we count oscillations at each epoch during training. The results demonstrate that QAT produces a significantly higher proportion of oscillating weights compared to $\lambda=0$. Furthermore, we observe that as we increase $\lambda$ a greater percentage of weights oscillates.
5 Experiments and Results
@@
In this section, we address the question: Is inducing weight oscillations during training sufficient to obtain the benefits of QAT? We answer this question affirmatively using empirical evidence based on the results of training ResNet-18 and Tiny Vit on the CIFAR-10 and Tiny-ImageNet datasets. This is both in the training-from-scratch (FS) setting and when fine-tuning (FT) pretrained models.
@@
In the following subsections we first describe the experimental setup, present accuracy results in FS and FT settings for models trained with different quantization levels for the quantizer in $\mathcal{R}_{\lambda}$ or QAT and finally, and report the cross-bit accuracy of the fine-tuned models. We train models at ternary (three possible values: -1, 0, 1), 3-bit and 4-bit. This is in line with contemporary research, where the emphasis lies on quantization at 4-bit and below, since the challenges of maintaining accuracy are more significant compared to quantization at higher bit widths.
5.1 Experimental Setup
@@
We conducted our experiments using the CIFAR-10 krizhevsky2009learning and Tiny-ImageNet Le2015TinyIV datasets. We evaluated three architectures; a multi-layer perceptron with five hidden layers and 256 neurons per layer (MLP5), ResNet-18 he2016deep and Tiny ViT wu2022tinyvit.
@@
For each architecture we used the Adam optimizer kingma2014adam and tested multiple configurations: a baseline model to establish optimal floating-point (FP) accuracy and PTQ performance, a model with QAT, and a model with regularization using Eq. (19). The two latter configurations are trained using 3-bit and 4-bit quantizers. In all our experiments we use the regularizer $\mathcal{R}_{\lambda}$ defined in Eq. (19) to induce oscillations (marked as “Oscillations” in the results table).
@@
Training-From-Scratch (FS): For the MLP5 architecture, we used a learning rate of $10^{-3}$ and regularization parameter $\lambda$=1. The ResNet-18 was trained with a learning rate of $10^{-3}$ and $\lambda$=0.75 (see Appx. A.2 for our hyperparameter selection). We modified the ResNet-18 architecture by replacing the input layer with a smaller $3\times 3$ kernel and adapting the final layer for 10-class classification of both ResNet-18 and Tiny ViT. Training proceeded for a maximum of 100 epochs with early stopping triggered after 10 epochs without improvement in validation performance. For quantized models, we monitored the quantized validation accuracy at the target bit precision, while for the baseline, we tracked floating-point accuracy.
@@
Fine-tuning (FT): We fine-tuned two ImageNet-1k deng2009imagenet pre-trained models on CIFAR-10 and Tiny-ImageNet: a Tiny ViT (learning rate: $10^{-4}$, $\lambda\in\{1,0.75,0.5\}$ depending on the bit) and a ResNet-18 (learning rate: $10^{-3}$, $\lambda\in\{1,0.75,0.5\}$ depending on the bit). To maintain compatibility with the pre-trained architectures, we up-sampled both CIFAR-10 and Tiny-Imagenet images to $224\times 224$ pixels. The $\lambda$ parameter selection process for Tiny ViT is detailed in Appendix A.2. Fine-tuning continued for up to 200 epochs on CIFAR-10 and 50 epochs for Tiny-ImageNet, with early stopping after 30 epochs without improvement, using the same accuracy metrics as training from scratch.
@@
Quantization: We implemented weight quantization using a per-tensor uniform symmetric quantizer as defined in Eq. (1). PTQ is applied in its most minimal form, by simply quantizing the weights without any calibration. QAT is used as defined in Eq.(5). The quantization range was determined by computing minimum and maximum values per layer. In our implementation of ResNet-18 (11M parameters) all layers except batch normalization were quantized, covering 99.96% of parameters. For Tiny ViT (5.5M parameters) quantization was applied to MLP, Self-Attention, and key-query-value projection layers, encompassing 97.18% of parameters. And lastly for the MLP5 model all layers were quantized. For Tiny-ImageNet models are trained at 3 and 4-bit precision only.
5.2 Results
The performance on the two datasets in training-from-scratch and fine-tuning settings is presented in the following sections, along with the observations about cross-bit generalization.
5.2.1 Performance on CIFAR-10
Training from Scratch:
@@ 5.2.1 Performance on CIFAR-10
Table 2 (A) shows the results from training an MLP and ResNet-18 from scratch on the CIFAR-10 dataset. Doing only regularization with Eq. (19) demonstrates improvements compared to the PTQ baseline. More importantly, it also matches the performance of QAT at bit widths of 3 and 4.
@@
For both models we see that at 3-bit and 4-bit, using the $\mathcal{R_{\lambda}}$ regularizer from Eq. (19) exhibits similar performance as QAT but with less variability. With both models, QAT and $\mathcal{R_{\lambda}}$ regularization are competitive with the full-precision baseline. Notably, both $\mathcal{R_{\lambda}}$ regularization and QAT significantly outperform PTQ when applied to the full-precision baseline.
Quantization MLP5 (FS) ResNet-18 (FS) Tiny ViT (FT) ResNet-18 (FT)
Baseline FP32 51.43 $\pm$ 0.39 83.26 $\pm$ 1.07 96.11 $\pm$ 0.31 88.50 $\pm$ 0.64
3-bit PTQ 20.97 $\pm$ 5.64 77.79 $\pm$ 4.00 11.56 $\pm$ 1.99 10.28 $\pm$ 0.48
3-bit QAT 50.53 $\pm$ 1.43 82.51 $\pm$ 2.14 88.13 $\pm$ 0.60 85.69 $\pm$ 1.83
3-bit Oscillations 48.48 $\pm$ 0.29 81.77 $\pm$ 0.46 88.68 $\pm$ 1.08 84.94 $\pm$ 1.59
4-bit PTQ 46.50 $\pm$ 0.76 82.11 $\pm$ 1.21 21.57 $\pm$ 5.33 35.56 $\pm$ 9.05
4-bit QAT 51.39 $\pm$ 0.60 82.66 $\pm$ 2.57 94.96 $\pm$ 0.33 87.71 $\pm$ 1.14
4-bit Oscillations 50.72 $\pm$ 0.47 83.74 $\pm$ 0.59 94.82 $\pm$ 0.51 87.08 $\pm$ 0.72
@@
Table 2: Performance comparison on CIFAR-10 dataset. Results show classification accuracy for MLP5, ResNet-18, and Tiny ViT across different quantization approaches and bit-widths. Models trained from scratch are marked FS, and fine-tuning experiments are marked FT. FT experiments are based on models pre-trained on ImageNet-1k. In all cases oscillations followed by quantization of the weights matches QAT accuracy. Results are means and standard deviations over 5 random seeds. PTQ results are from the FP32 baseline.
Fine-tuning:
@@
Table 2 (B) summarizes the test accuracies for fine-tuning using our $\mathcal{R_{\lambda}}$ regularization and QAT on ResNet-18 and Tiny ViT architectures with CIFAR-10 and Tiny-ImageNet. The observations are roughly in line with the results observed for training from scratch in the previous section.
@@
For CIFAR-10 as with the case for training from scratch, with both ResNet-18 and Tiny Vit, we see an increase in performance compared to PTQ when using the regularization in Eq. (19). Regularization with $\mathcal{R_{\lambda}}$ and QAT show comparable performance when quantized at 3 bits and 4 bits, while achieving test accuracy close to the full-precision model at 4-bits.
Performance comparison and related discussions for ternary quantization are presented in Appendix A.3.
Robustness to Cross-bit Quantization:
@@
[…]